OpenStudy (anonymous):

p-5-13p=150-11 need a solution

OpenStudy (anonymous):

-12p = 150-11+5 -12p = 150-6 -12p = 144 p = 12

jimthompson5910 (jim_thompson5910):

p-5-13p=150-11 -12p-5=150-11 -12p-5=139 -12p=139+5 -12p=144 p=(144)/(-12) p=-12

OpenStudy (anonymous):

one says negative and one says positive which one is it

jimthompson5910 (jim_thompson5910):

p = -12 is the right answer, integral made a typo. Notice how he has -12p = 144, but the negative is missing.

OpenStudy (anonymous):

woops sorry u're right jim

OpenStudy (anonymous):

thanks for clearing that up